OpenStudy (anonymous):

Runner A is initially 2.6 km west of a flagpole and is running with a constant velocity of 3.8 km/h due east. Runner B is initially 7.0 km east of the flagpole and is running with a constant velocity of 7.8 km/h due west. What will be the distance of the two runners from the flagpole when their paths cross? km from the flagpole due (east)

OpenStudy (anonymous):

i dont hve to convert anything its all in km and km/h right

OpenStudy (shane_b):

Right...let me get a piece of paper and draw this one :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

its like the car but instead of a pole they use a flag pole as a point of reference

OpenStudy (shane_b):

Ok...one guy is running east at 3.8km/h and the other is running west at 7.8km/h. Their combined rate of closure is 7.8+3.8=11.6km/hr and they have to cover a total distance of 7+2.6km=9.6km. \[Speed = \frac{distance}{time}\]\[11.6km/h=\frac{9.6km}{t}\]First solve for t. Then use t x one runner's speed to determine how far that runner got in that amount of time. That should tell you where they meet.

OpenStudy (anonymous):

what is tx?

OpenStudy (shane_b):

distance = speed x time so it's: time * one runner's speed

OpenStudy (anonymous):

9.6/3.8?

OpenStudy (shane_b):

\[11.6km/h=\frac{9.6km}{t}\]\[t=\frac{9.6km}{11.6km/h}\]

OpenStudy (shane_b):

(and note that that will be in hours)

OpenStudy (anonymous):

0.827 km?

OpenStudy (shane_b):

0.83 hours

OpenStudy (anonymous):

but they ask for km in the answers

OpenStudy (shane_b):

We're not done yet :)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

what next

OpenStudy (shane_b):

At this point you know how long it took them to cover the entire distance given their combined speeds. Now all you need to do is figure out where one of the runners will be at that time: Let's go with the runner heading east at 3.8 km/h:\[distance=speed*time=(3.8km/h)(0.83h)=?\]

OpenStudy (anonymous):

souldn't u only round the final answer or it may be wrong?

OpenStudy (shane_b):

It probably won't matter in this problem but you can round it at the end if you like.

OpenStudy (anonymous):

now plug in 7.0 for the next runner then subtract the two numbers?

OpenStudy (shane_b):

Not exactly. Here's how I would do it...using 5 significant digits for the time should be sufficient: Runner 1 @ 3.8 km/h:\[(3.8km/h)(0.8276h)=3.1449km\] Runner 2 @ 7.8 km/h:\[(7.8km/h)(0.8276h)=6.4553km\] That looks right...the distance was about 9.6km total if you remember from the beginning.

OpenStudy (anonymous):

now subtract those two?

OpenStudy (shane_b):

So the guy running right went 3.1449km and the flagpole was 3.8km from him. 3.8km-3.1449km = 0.6851km to the left of the flag pole when the runners meet.

OpenStudy (shane_b):

You could also do the reverse using the other runner's speed and the time but the answer would be the same.

OpenStudy (anonymous):

pellet that was my last submission in webassign so all this effert gone to waist

OpenStudy (shane_b):

yuck. Well, hopefully you learned a little something anyway

OpenStudy (anonymous):

yeah i got a easy one comin next so look for it

OpenStudy (shane_b):

ok